OpenStudy (anonymous):

Consider the curve defined by x^2 + xy + y^2=27

OpenStudy (anonymous):

a.) Write an expression for the slope of the curve at any point (x,y) b.) Determine whether the lines tangent to the curve at the x-intercepts of the curve are parallel. Show the analysis that leads to your conclusion c.) Find the points on the curve where the lines tangent to the curve are vertical

OpenStudy (anonymous):

a) Find dy/dx via implicit differentiation. Do you know how to do that?

OpenStudy (anonymous):

yeah i got -2x-y/(x+2y) is that correct and the answer to part a.)? @SithsAndGiggles

OpenStudy (anonymous):

Yes, that's right. Now find the x-intercepts.

OpenStudy (anonymous):

how would i find the x-intercepts

OpenStudy (anonymous):

set y=0

OpenStudy (anonymous):

to the what i got from doing dy/dx -2x-y/(x+2y) = 0 ?

OpenStudy (anonymous):

Using the original curve, \[x^2+xy+y^2=27,\] and setting y = 0 (as @Hoa suggested, since all x-intercepts have the ordered pair (x, 0) for some x), you have \[x^2+x(0)+0^2=27\\ x^2=27\]